# SynPlay: Importing Real-world Diversity for a Synthetic Human Dataset

### v1

@@ 3.1 Other design factors
Characters. We have designed 456 virtual characters using the Character Creator [47], where each character is involved in multiple game scenarios. To vary the appearance of the characters and avoid generating biases, each character was uniquely designed with gender, skin color, age, height, obesity (body type), hair (styles and colors), and outfits. We kept the gender ratio between male and female at 1:1 and the ratio of skin color among white, black, yellow, and brown at 1:1:1:1. For age, each character was designed to fall into one of three categories: child, middle-aged, and elderly, and the ratio was set at 1:2:1. For each gender and age group, heights were modeled to follow a bell-shaped distribution, resulting in an overall dataset range of 140 to 190 cm. We manually designated every character with a unique outfit, while setting the hair and obesity aspect as diverse as possible.
Backgrounds. For each scenario, we set different environmental factors: sites, lighting conditions, and weather. A game takes place at one of ten custom-built sites. These include five urban locations (three common city areas, a construction site, and a factory site) and five natural sites (a green area, a snowy field, a desert, a meadow, and a beach). Multiple locations within each site map can be used as local playgrounds. To vary the lighting conditions, we take into account five different times of day: dawn, morning, noon, afternoon, and sunset. We consider three types of weather: sunny, foggy, and rainy. All of these are randomly determined for each scenario.
@@ 4.1 General tasks: detection and segmentation
The main tasks are conducted on aerial-view datasets, which feature a wider range of human appearances, making them ideal for validating the design philosophy behind the SynPlay dataset. We also conduct experiments on ground-view datasets to evaluate SynPlay on a more widely studied task in the community.
Table 1: Comparison with other synthetic datasets on aerial-view human detection and semantic segmentation. The numbers in parentheses are the gaps from the model trained without synthetic data (‘real’). Positive and negative gaps are indicated in green and red fonts, respectively. The best accuracy for each setting is shown in bold. Notations: ‘+ real’ represents a model pre-trained with synthetic data and fine-tuned on a ‘real’ dataset, where ‘real’ is a training set of dataset used for evaluation. ‘s’, ‘m’, and ‘l’ represent three YOLO v8 models with different architectures. (a) human detection data in training VisDrone [54] Okutama-action [5] Semantic Drone [1] s m l s m l s m l   real 19.72/47.43 21.14/49.52 21.60/51.10 27.40/75.17 28.99/76.60 31.53/78.78   44.00/  77.20   44.52/  78.52   42.62/  79.87 Archangel [45]   0.23/  0.63   0.38/  0.98   0.59/  1.48   2.59/  8.45   3.90/10.13   2.83/  9.12     0.64/    1.59     2.42/    5.37     0.94/    1.62 SynDrone [40]   0.31/  0.81   0.36/  0.84   0.71/  1.89   0.00/  0.00   0.00/  0.01   0.00/  0.00     0.00/    0.00     0.00/    0.00     0.00/    0.00 SynPlay   5.29/11.75   4.31/  9.12   2.79/  5.87 12.74/40.86   8.19/25.43   8.15/25.23     7.02/  12.21     9.60/  15.51   15.71/  23.59 Archangel + real 18.77/45.39 20.25/48.52 20.82/49.51 30.72/80.35 32.36/80.63 31.71/79.63   46.60/  74.07   48.60/  75.86   44.62/  73.23 (-0.95/-2.04) (-0.89/-1.00) (-0.78/-1.59) (+3.32/+5.18) (+3.37/+4.03) (+0.18/+0.85) (  +2.60/  -2.13) (  +4.08/  -2.66) (  +1.00/  -6.64) SynDrone + real 18.78/45.79 20.94/49.44 21.97/51.51 29.70/77.71 31.39/79.42 31.24/78.71   50.93/  82.28   53.71/  85.47   59.59/  85.02 (-0.94/-1.64) (-0.20/-0.08) (+0.37/+1.41) (+2.30/+2.54) (+2.40/+2.82) (-0.29/-0.07) (  +6.93/  +5.08) (  +9.19/  +6.95) (+16.97/  +5.15) SynPlay + real 20.88/49.31 22.34/52.12 22.98/52.93 32.47/81.60 31.96/81.13 33.17/82.52   66.52/  90.33   69.46/  91.35   68.82/  91.37 (+1.16/+1.88) (+1.20/+2.60) (+1.38/+1.83) (+5.07/+6.43) (+2.97/+4.53) (+1.64/+3.74) (+22.52/+13.13) (+24.94/+12.83) (+26.20/+11.50) (b) semantic seg. Semantic Drone Aeroscapes [1] [35]       0.66 22.25     0.74   0.04     0.07   0.00     8.03   6.44     9.28 20.61 (  +8.62) (-1.64)     5.56 24.59 (  +4.90) (+2.34)   23.32 32.19 (+22.66) (+9.94)
Aerial-view tasks. Table 1 shows the results for aerial-view human detection and semantic segmentation tasks. Overall, for both tasks, using SynPlay in training provides remarkably better accuracy than all the compared cases, including real and all the other variations involving other synthetic data.
@@
Table 2: Impact of SynPlay on MS COCO (person category). Notation: ‘SynPlay-UGV’ and ‘SynPlay-all’ are a UGV subset of SynPlay and the entire SynPlay, respectively. (a) human detection data in training s m l   real 46.19/65.91 50.10/69.86 52.52/72.15 SynPlay-UGV + real 46.53/66.18 50.70/70.37 52.69/72.29 (+0.34/+0.27) (+0.60/+0.51) (+0.17/+0.14) SynPlay-all + real 46.84/66.70 51.12/70.74 53.00/72.59 (+0.65/+0.79) (+1.02/+0.88) (+0.48/+0.44) (b) sem.seg.   15.10 20.18 (+5.08) 21.57 (+6.47)    Table 3: Synergy impact with MS COCO on aerial-view human detection. YOLO v8 model with a medium size architecture is used. data in training VisDrone Okutama-action Semantic Drone   real 21.14/49.52 28.99/76.60 44.52/78.52 COCO   7.16/16.46 15.17/48.28 34.74/56.39 SynPlay   4.31/  9.12   8.19/25.43   9.61/15.52 COCO + SynPlay 11.49/25.20 14.68/49.82 18.60/31.03 COCO + real 22.11/51.73 32.26/80.10 65.72/89.20 SynPlay + real 22.34/52.13 31.96/81.13 69.46/91.35 COCO + SynPlay + real 22.78/53.01 33.82/79.44 73.52/92.80
Ground-view tasks. Table 3 explores the impact of using SynPlay for the general computer vision tasks of ground-view human detection and semantic segmentation. We also evaluate how models perform when only the subset with matching viewpoint (i.e., UGV images in SynPlay) is used in training. Overall, using the entire SynPlay yields the highest accuracy on both tasks, while using the UGV-subset still outperforms the model trained without SynPlay. These results demonstrate that our insight in ensuring diversity by varying the camera viewpoints is effective even in tasks that do not contain such multiple viewpoints. In addition, the greater improvement in semantic segmentation over object detection shows that ensuring diversity is more effective in tasks that require more detailed human representation models.
@@ 4.2 Data-scarce tasks: Few-shot and cross-domain learning
In this section, we compare SynPlay with other synthetic datasets on its ability to meet the demand for additional data in data-scarce tasks. For data-scarce tasks, we adopt few-shot and cross-domain learning tasks on aerial-view human detection, which suffers more severely from a lack of training data than ground-view detection. Following the data-scarce task setups of [43], we train models with two few-shot regimes using 20 and 50 images of VisDrone (denoted by ‘Vis-20/50’). We test the models on VisDrone, Okutama-action, and Semantic Drone where the evaluations done on the last two datasets can be seen as ‘cross-domain’. To attenuate the potential random effects that may arise when selecting real training images, all reported numbers are average accuracy over three runs.
As baseline methods leveraging synthetic data in training, we use a pretrain-finetune strategy (PT-FT) and Progressive Transformation Learning (PTL) [43]. PTL is a progressive data augmentation approach that iteratively expands the training set by adding a subset of synthetic data, which is transformed to look real. In each PTL iteration, a subset of the synthetic data is selected, such that synthetic data that is closer to the real dataset is selected more often. For the data-scarce tasks experimented in [43], PTL was better than PT-FT while both outperformed the cases without synthetic data. We used RetinaNet [28] as the detector. [^5] PTL was designed to be suitable for RetinaNet. For a fair comparison between PTL and PT-FT, we used RetinaNet instead of YOLO v8 for this experiment.
[…]
@@
Comparison with other synthetic data. In Table 4, we compare the detection accuracy of the models trained with different synthetic datasets on the few-shot and cross-domain learning tasks. With PT-FT, SynPlay achieved significantly better accuracy than other synthetic datasets across all three datasets. With PTL, SynPlay performed the best on VisDrone and Okutama-action. These results were consistent for both Vis-20 and Vis-50 settings. Even on Semantic Drone, which shows an unusual performance trend, the best performance was achieved when SynPlay was used via PT-FT.
@@ 4.3 Image quality evaluation
In Table 5, for all training datasets involved in our experiments, we calculate FID (Fréchet Inception Distance) [23] to assess their fidelity and diversity. SynPlay presents the best score compared to other synthetic datasets, which is a result that aligns well with our task results. These results suggest that SynPlay’s superior task performance is achieved by better fidelity and diversity, which are our goals in designing SynPlay. Moreover, the better FID of SynPlay compared to MS COCO, which mainly includes ground-view images, is also reported, supporting the hypothesis that adopting multiple viewpoints effectively diversifies human appearances.
@@ 5 Discussion and Conclusion
Table 6: Proportion of nadir-view instances in synthetic datasets used in data-scarce tasks. Instances with camera viewing angle from ground greater than 71.57 are considered as nadir-view instances. [^7] This criterion is the largest viewing angle from ground on Archangel*.
@@
  • The performance gain acquired by incorporating a synthetic data during training is remarkably large for Semantic Drone when compared to the other two cases, while SynPlay’s incorporation showcasing the most substantial gain.
Most human instances in Semantic Drone are taken from nadir views, while VisDrone and Okutama-action have instances captured with fewer nadir-views. The portion of nadir-view instances in SynPlay is the smallest among the synthetic datasets used in data-scarce tasks (Table 7). As PTL continues to prioritize samples from synthetic data that closely resemble the seed data (i.e., VisDrone) for training, the reduced selection of nadir-view instances from the SynPlay may result in a lower gain (first phenomenon). On the other hand, the second phenomenon indicates that ensuring greater diversity via using supplemental synthetic data has greater impact on Semantic Drone, which lacks diversity due to its limited viewpoints. Moreover, SynPlay that is less similar to Semantic Drone while being more diverse in relation to the compared synthetic datasets, shows the largest impact, supporting our claim that improving the diversity is generally effective in constructing a better synthetic data.
Broader impact. Utilizing real human datasets frequently entails inherent privacy concerns. We hope that our endeavors to enhance synthetic human data, moving it one step closer to real-world fidelity, will contribute to alleviating these challenges.
Limitations. SynPlay was developed to provide richer representations of human appearance for tasks that involve localizing human in the scenes. We recognize the significance of incorporating distinctive features from diverse object categories. For future work, we are aiming to expand the SynPlay to encompass a wider array of categories, thereby enriching its training capabilities.
Conclusion. What motion a human performs and where a person is viewed from are two crucial factors that make a difference in how a human looks. We create a synthetic human dataset called SynPlay with the aim of expanding the diversity human appearance by diversifying these factor. Ensuring the diversity allowed SynPlay to have a greater positive impact towards model training when compared to other counterparts (train-from-scratch, using other synthetic data) on aerial-view/ground-view object detection and semantic segmentation. This positive impact of SynPlay becomes even greater in data-scarce tasks, where synthetic data is strongly desired as supplemental training data.
Acknowledgement. This research was sponsored by the Defense Threat Reduction Agency (DTRA) and the DEVCOM Army Research Laboratory (ARL) under Grant No. W911NF2120076. This research was also sponsored in part by the Army Research Office and Army Research Laboratory (ARL) under Grant Number W911NF-21-1-0258. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office, Army Research Laboratory (ARL) or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@
Appendix A Comparison to Other Datasets
  1. 1.
  2. 2.
[…]
  3. 3.
  4. 4.
    Rule-guided design can utilize significantly larger range of human motions compared to detail-guided design.
[…]
Moreover, ${}^{\text{th}}$ observation supports that our proposed rule-guided design is successful in securing the diversity of human motions in the set. It is noteworthy that while SURREAL [50] (constructed with ‘detail+mocap’) contains a comparable number (6.5M) of human instances as SynPlay, the number of motions manifested in the dataset is extremely limited when compared to SynPlay (23 vs. uncountable).
@@ Appendix A Comparison to Other Datasets
VOC 12 [17] real 10K 11.5K 2.48 daily uncountable near
KITTI [18] real 4.5K 7.5K 2.52 daily 2 near
COCO Dev17 [29] real 649K 164K 9.72 daily uncountable near
MPII Human Pose [3] real 40K 24.9K 1.61 daily 20 near
Cityscapes [12] real 21.4K 5K 7.85 daily 2 near
ADE20K [53] real 30K 27.5K 4.36 daily uncountable near
Human-Art [26] real 123K 50K 2.46 art uncountable near
@@
SURREAL [50] synth 6.5M 6.5M 1 detail+mocap 23 near
[…]
UnrealPerson [52] synth 120K 120K 1 2 near
[…]
aerial-view
Okutama-action [5] real $\cdot$ 77K $\sim$9 detail 12 med
Semantic Drone [1] real 1.5K 400 4.16 daily unspecified med
[…]
VisDrone [54] real 109K 40.0K 15.42 daily unspecified med

### v2

@@ 3.1 Other design factors
Characters. We designed 456 virtual characters using the Character Creator, with each character participating in multiple game scenarios. To vary appearances and avoid biases, each character was uniquely designed by gender, skin color, age, height, obesity (body type), hair (styles and colors), and outfits. We kept the gender ratio between male and female at 1:1 and the ratio of skin color among white, black, yellow, and brown at 1:1:1:1. For age, each character was designed to fall into one of three categories: child, middle-aged, and elderly, and the ratio was set at 1:2:1. For each gender and age group, heights were modeled to follow a bell-shaped distribution, resulting in an overall dataset range of 140 to 190 cm. We manually designated every character with a unique outfit, while setting the hair and obesity aspect as diverse as possible.
Backgrounds. For each scenario, we set different environmental factors: sites, lighting conditions, and weather. A game takes place at one of ten custom-built sites, including five urban locations (three city areas, a construction site, and a factory) and five natural sites (a green area, snowy field, desert, meadow, and beach). Multiple locations within each site map can serve as playgrounds. To vary lighting, we consider five different times of day: dawn, morning, noon, afternoon, and sunset. We use three types of weather: sunny, foggy, and rainy. All of these are randomly determined for each scenario.
@@ 4 Task Evaluation
Table 1: Comparison with other synthetic datasets on aerial-view human detection and semantic segmentation. The numbers in parentheses are the performance gaps from the model trained without synthetic data (‘real’). Positive and negative performance gaps are indicated in green and red fonts, respectively. The best accuracy for each setting is shown in bold. Notations: ‘+ real’ represents a model pre-trained with synthetic data and fine-tuned on the train split of the corresponding ‘real’ dataset. Performances are evaluated on test split of corresponding ‘real’ data. ‘s’, ‘m’, and ‘l’ represent three YOLO v8 models with different architectures.
(a) human detection
data in training VisDrone [PZhuTPAMI2022] Okutama-action [MBarekatainCVPRW2017] Semantic Drone [ICGlink]
s m l s m l s m l
  real 19.72/47.43 21.14/49.52 21.60/51.10 27.40/75.17 28.99/76.60 31.53/78.78   44.00/  77.20   44.52/  78.52   42.62/  79.87
Archangel [YShenAccess2023]   0.23/  0.63   0.38/  0.98   0.59/  1.48   2.59/  8.45   3.90/10.13   2.83/  9.12     0.64/    1.59     2.42/    5.37     0.94/    1.62
SynDrone [GRizzoliICCVW2023]   0.31/  0.81   0.36/  0.84   0.71/  1.89   0.00/  0.00   0.00/  0.01   0.00/  0.00     0.00/    0.00     0.00/    0.00     0.00/    0.00
SynPlay   5.29/11.75   4.31/  9.12   2.79/  5.87 12.74/40.86   8.19/25.43   8.15/25.23     7.02/  12.21     9.60/  15.51   15.71/  23.59
Archangel + real 18.77/45.39 20.25/48.52 20.82/49.51 30.72/80.35 32.36/80.63 31.71/79.63   46.60/  74.07   48.60/  75.86   44.62/  73.23
(-0.95/-2.04) (-0.89/-1.00) (-0.78/-1.59) (+3.32/+5.18) (+3.37/+4.03) (+0.18/+0.85) (  +2.60/  -2.13) (  +4.08/  -2.66) (  +1.00/  -6.64)
SynDrone + real 18.78/45.79 20.94/49.44 21.97/51.51 29.70/77.71 31.39/79.42 31.24/78.71   50.93/  82.28   53.71/  85.47   59.59/  85.02
(-0.94/-1.64) (-0.20/-0.08) (+0.37/+1.41) (+2.30/+2.54) (+2.40/+2.82) (-0.29/-0.07) (  +6.93/  +5.08) (  +9.19/  +6.95) (+16.97/  +5.15)
SynPlay + real 20.88/49.31 22.34/52.12 22.98/52.93 32.47/81.60 31.96/81.13 33.17/82.52   66.52/  90.33   69.46/  91.35   68.82/  91.37
(+1.16/+1.88) (+1.20/+2.60) (+1.38/+1.83) (+5.07/+6.43) (+2.97/+4.53) (+1.64/+3.74) (+22.52/+13.13) (+24.94/+12.83) (+26.20/+11.50)
(b) semantic seg.
Semantic Drone Aeroscapes [ICGlink] [INigamWACV2018]       0.66 22.25     0.74   0.04     0.07   0.00     8.03   6.44     9.28 20.61 (  +8.62) (-1.64)     5.56 24.59 (  +4.90) (+2.34)   23.32 32.19 (+22.66) (+9.94)
@@ 4.1 General tasks: detection and segmentation
The main tasks use aerial-view datasets, featuring a wider range of human appearances, making them ideal for validating the design philosophy of the SynPlay. We also conduct experiments on ground-view datasets to evaluate SynPlay on a more widely studied task in the community.
Aerial-view tasks. Table 1 shows the results for aerial-view human detection and semantic segmentation tasks. Overall, for both tasks, using SynPlay in training provides remarkably better accuracy than all the compared cases, including real and all other variations involving other synthetic data.
@@
Ground-view tasks. Table 2 explores the impact of using SynPlay for the general computer vision tasks of ground-view human detection and semantic segmentation. We also evaluate how models perform when only the subset with matching viewpoint (i.e., UGV images in SynPlay) is used in training. Overall, using the entire SynPlay yields the highest accuracy on both tasks, while using the UGV-subset still outperforms the model trained without SynPlay. These results demonstrate that our insight in ensuring diversity by varying the camera viewpoints is effective even in tasks that do not contain such multiple viewpoints. In addition, the greater improvement in semantic segmentation over object detection shows that ensuring diversity is more effective in tasks that require more detailed human representations.
Table 2: Impact of SynPlay on MS COCO (person category). Notation: ‘SynPlay-UGV’ and ‘SynPlay-all’ are a UGV subset of SynPlay and the entire SynPlay, respectively.
(a) human detection
data in training s m l
  real 46.19/65.91 50.10/69.86 52.52/72.15
SynPlay-UGV + real 46.53/66.18 50.70/70.37 52.69/72.29
(+0.34/+0.27) (+0.60/+0.51) (+0.17/+0.14)
SynPlay-all + real 46.84/66.70 51.12/70.74 53.00/72.59
(+0.65/+0.79) (+1.02/+0.88) (+0.48/+0.44)
(b) sem.seg.
  15.10
20.18
(+5.08)
21.57
(+6.47)
@@
Table 3: Synergy impact with MS COCO on aerial-view human detection. YOLO v8 model with a medium size architecture is used.
data in training VisDrone Okutama-action Semantic Drone
  real 21.14/49.52 28.99/76.60 44.52/78.52
COCO   7.16/16.46 15.17/48.28 34.74/56.39
SynPlay   4.31/  9.12   8.19/25.43   9.61/15.52
COCO + SynPlay 11.49/25.20 14.68/49.82 18.60/31.03
COCO + real 22.11/51.73 32.26/80.10 65.72/89.20
SynPlay + real 22.34/52.13 31.96/81.13 69.46/91.35
COCO + SynPlay + real 22.78/53.01 33.82/79.44 73.52/92.80
@@ 4.2 Data-scarce tasks: Few-shot and cross-domain learning
We compare SynPlay with other synthetic datasets in addressing data scarcity in aerial-view human detection, which lacks training data more severely than ground-view detection. Following the data-scarce task setups of  [YShenCVPR2023], we evaluate few-shot and cross-domain learning. Models are trained with 20 and 50 VisDrone images (‘Vis-20/50’) and tested on VisDrone, Okutama-action, and Semantic Drone, with the latter two serving as cross-domain evaluations. To mitigate randomness in selecting real training images, all reported accuracy values are averaged over three runs.
As baseline methods leveraging synthetic data in training, we use a pretrain-finetune strategy (PT-FT) and Progressive Transformation Learning (PTL) [YShenCVPR2023]. PTL is a progressive data augmentation approach that iteratively expands the training set by adding a subset of synthetic data, which is transformed to look real. In each PTL iteration, a subset of the synthetic data is selected, such that synthetic data that is closer to the real dataset is selected more often. For the data-scarce tasks experimented in [YShenCVPR2023], PTL was better than PT-FT while both outperformed the cases without synthetic data. We used RetinaNet [TLinICCV2017] as the detector. [^4] PTL was designed to be suitable for RetinaNet. For a fair comparison between PTL and PT-FT, we used RetinaNet instead of YOLO v8 for this experiment.
Table 4: Few-shot and cross-domain learning accuracy on aerial-view human detection. To clarify, the accuracy on Okutama-action and Semantic Drone refers to cross-domain learning performance. Notation: ‘Archangel*’ is an expanded ‘Archangel’ to be pose-diversified [YShenArXiv2024].
@@
Comparison with other synthetic data. In Table 4, we compare the detection accuracy of the models trained with different synthetic datasets on the few-shot and cross-domain learning tasks. With PT-FT, SynPlay achieved significantly better accuracy than other synthetic datasets across all three datasets. With PTL, SynPlay performed the best on VisDrone and Okutama-action for both Vis-20 and Vis-50 settings. Even on Semantic Drone, which shows an unusual performance trend, the best performance was achieved when SynPlay was used via PT-FT.
@@ 4.3 Image quality evaluation
In Table 5, for all synthetic data used in training, we calculate FID (Fréchet Inception Distance) [MHeuselNeurIPS2017] to assess their fidelity and diversity. SynPlay presents the best score, which aligns well with our task results. It suggests that SynPlay’s superior task performance is achieved by better fidelity and diversity, which are our goals in designing SynPlay. Moreover, the better FID of SynPlay compared to MS COCO, which mainly includes ground-view images, is also reported, supporting the hypothesis that adopting multiple viewpoints effectively diversifies human appearances.
@@ 5 Discussion and Conclusion
Table 6: Proportion of nadir-view instances in synthetic datasets used in data-scarce tasks. Instances with camera viewing angle from ground greater than 71.57 (largest viewing angle from ground in Archangel*) are considered as nadir-view instances.
@@
  • The performance gain acquired by incorporating a synthetic data during training is remarkably large for Semantic Drone when compared to other two cases, while SynPlay showcasing the most substantial gain.
Most human instances in Semantic Drone are taken from nadir views, while VisDrone and Okutama-action have fewer nadir-views. The portion of nadir-view instances in SynPlay is the smallest among the synthetic datasets used in data-scarce tasks (Table 6). As PTL continues to prioritize samples from synthetic data that closely resemble the seed data (i.e., VisDrone) for training, the reduced selection of nadir-view instances from the SynPlay may result in a lower gain (first phenomenon). On the other hand, the second phenomenon indicates that ensuring greater diversity via using supplemental synthetic data has greater impact on Semantic Drone, which lacks diversity due to its limited viewpoints. Moreover, SynPlay that is less similar to Semantic Drone while being more diverse than other synthetic datasets, shows the largest impact, supporting our claim that improving the diversity is effective in constructing a better synthetic data.
Conclusion. We presented SynPlay, a large-scale synthetic human dataset that addresses the critical challenge of identifying humans in aerial views where subjects appear as only tens of pixels. SynPlay departs from traditional datasets by introducing a rule-guided motion generation, enabling effectively uncountable motion variations and spontaneous interactions. Through a multi-perspective setup, SynPlay uniquely supports near-to-far human localization across extreme viewpoint variations. Extensive experiments confirm that SynPlay significantly improves human detection, segmentation, and keypoint estimation, particularly in few-shot and cross-domain settings. By advancing both dataset design and practical model performance, SynPlay establishes a new foundation for long-range human analysis and aerial-view perception.
@@
Appendix A Comparison to Other Human Datasets
@@
VOC 12 [MEveringhamIJCV2015] real 10K 11.5K 2.48 daily uncountable near
KITTI [AGeigerCVPR2012] real 4.5K 7.5K 2.52 daily 2 near
COCO Dev17 [TLinECCV2014] real 649K 164K 9.72 daily uncountable near
MPII Human Pose [MAndrilukaCVPR2014] real 40K 24.9K 1.61 daily 20 near
Cityscapes [MCordtsCVPR2016] real 21.4K 5K 7.85 daily 2 near
ADE20K [BZhouCVPR2017] real 30K 27.5K 4.36 daily uncountable near
Human-Art [XJuCVPR2023] real 123K 50K 2.46 art uncountable near
GTA5 [SRichterECCV2016] synth 1.4M 1.4M 1 $20K_{\text{pose}}$ near
SURREAL [GVarolCVPR2017] synth 6.5M 6.5M 1 detail+mocap 23 near
SOMAset [IBarbosaCVIU2018] synth 100K 100K 1 detail+mocap $250_{\text{pose}}$ near
PersonX [XSunCVPR2019] synth 273K 273K 1 $4_{\text{pose}}$ near
UnrealPerson [TZhangCVPR2021] synth 120K 120K 1 2 near
AGORA [PPatelCVPR2021] synth $\cdot$ 19K $\sim1$5 detail+mocap $4,240_{\text{pose}}$ near
BEDLAM [MBlackCVPR2023] synth $\cdot$ 380K $\sim1$0 detail+mocap $2,311_{\text{pose}}$ near
aerial-view
Okutama-action [MBarekatainCVPRW2017] real $\cdot$ 77K $\sim$9 detail 12 med
Semantic Drone [ICGlink] real 1.5K 400 4.16 daily unspecified med
@@
VisDrone [PZhuTPAMI2022] real 109K 40.0K 15.42 daily unspecified med
@@
  1. 1.
  2. 2.
[…]
  3. 3.
  4. 4.
    Rule-guided design, which is only leveraged in the SynPlay, can utilize significantly larger range of human motions compared to detail-guided design.
[…]